collincaroll44:

A radio active isotope decays according to the exponential decay equation where t is in days. Round to the thousandths place. For the half life: The half life is the solution (t) of the equation : a/2=ae^−3.568t

sillybilly123:

\(\dfrac{a}{2}=a e^{ - 3.568 ~ T}\) \(\implies \frac{1}{2}=e^{ - 3.568 ~ T}\) \( \implies \ln ( \frac{1}{2} )= \ln e^{ - 3.568t ~ T }\) \(\implies \ln (1) - \ln (2) = - 3.568 ~ T\) finish that off :) NB, using *T* rather than *t*, as this is a specific instance of the decay equation. ie the \(\dfrac{a}{2}=a e^{ - 3.568 ~ t}\) is only true for t = the half-life = T.